OpenStudy (anonymous):

If the angle measures of the interior angles of a pentagon are 130 degrees, 94 degrees, 109 degrees, 112 degrees, and X degrees. Find X?!

OpenStudy (anonymous):

My question is from the polygon angle sum therom btw.

Directrix (directrix):

Yes, it is. The sum of the interior angle measures is (n-2)*180. n = 5 for the 5 sides of a pentagon. So, first up, what is the sum of the interior angle measures of a pentagon? @XtraStupid

Directrix (directrix):

To get x, we need that sum. 130 + 94 + 109 + 112 + x = ? --> the sum goes here.

Directrix (directrix):

(n-2)*180. n = 5 (5-2)*180 = ? @XtraStupid

OpenStudy (anonymous):

These are the only answers though. X= 95, x=105, x=108, or x=275 ??

OpenStudy (anonymous):

I still dont know what the answer is :(

Directrix (directrix):

If you would work with me and not look at the answers, that would be great. We cannot solve this problem until you find the sum of the angles measures of a pentagon. So, for the third time, I am asking you to perform this calculation (5-2)*180 = ? In short, what is three times 180 equal to? Please post the answer. We need it for the equation. We are nowhere near the answer so don't look at the options.

Directrix (directrix):

What is three times 180 equal to? Please post the answer.

OpenStudy (anonymous):

Oh my bad. It's 540

Directrix (directrix):

Okay. 130 + 94 + 109 + 112 + x = 540 ^^ Solve that equation for x. Post what you get so we can compare answers, okay?

OpenStudy (anonymous):

x would equal 95 :D omgeeee i think i did it lol!!!

Directrix (directrix):

95 is what I got, too.

OpenStudy (anonymous):

100 percent. you helped so much thank you (:

Directrix (directrix):

You are welcome.